# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ A.2 DiD setups with continuous or multi-valued treatments
Our paper focuses on binary treatments, but many treatments take multiple values or are even continuous. A number of recent papers have studied this particular type of treatment design; see, e.g., Callaway2021b; CallawayGoodmanBaconSantAnna2024 and deChaisemartin2022-continuous-treatment; deChaisemartinDHaultfoeuilleVazquezBare2024. Here we focus on a two-period setting in which no unit is treated in period one and some units receive a treatment with varying intensities (or doses) in period two. Most of the key results that distinguish multi-valued from binary treatments are evident with two periods (CallawayGoodmanBaconSantAnna2024).
We now need to define potential outcomes that reflect varying treatment intensity. We denote $Y_{i,t}(0,d)$ as potential outcomes for unit $i$ in period $t$ if they are untreated in period one and receive treatment dosage $d$ in period two. As we focus on setups where all units are untreated in period one, we simplify notation and index potential outcomes by treatment intensity in period two, i.e., $Y_{it}(d)=Y_{i,t}(0,d)$. An important feature is that $d$ is not restricted to $\{0,1\}$ and can take on richer treatment intensities instead. We denote the treatment dosage for unit $i$ as $D_{i}$ in period two and stress that, in this context, our notion of the treatment group is tied to units’ treatment dosage, i.e., groups are defined by their treatment dosage in period 2.
@@
reflect the average effect of dose $d$ relative to no treatment. $ATT(d|d^{\prime})$ is the average treatment effect for units that experienced dose $d^{\prime}$; when $d^{\prime}=d$ it is the $ATT$ among units that received dose $d$. $ATE(d)$ is defined analogously, except that it is the effect on the overall population. Of course, one can also aggregate these dose-specific parameters to form more precisely estimable summary quantities; see, e.g., Callaway2021b.
$ATT(d|d)$ and $ATE(d)$ provide average treatment effects in levels and so one reason why they vary could be because $d$ itself varies. To account for the differences in $d$, one may be interested in “per-dosage” effects
@@
$ACRT(d|d)$ equals the derivative of the average potential outcome in period two for units that received dose $d$ evaluated at $d$—this is equivalent to the derivative of $ATT(l|d)$ with respect to $l$, evaluated at $l=d$. $ACR(d)$ can be interpreted analogously. [^35] For discrete treatments, ACR’s are defined in a similar way but with a slightly different notation to accommodate the discreteness of $d$, i.e., $ACRT(d_{j}|d_{k})=\mathop{}\!\mathbb{E}[Y_{t=2}(d_{j})-Y_{t=2}(d_{j-1})|D=d_{k}]$, and $ACR(d_{j})=\mathop{}\!\mathbb{E}[Y_{t=2}(d_{j})-Y_{t=2}(d_{j-1})]$.
The relevant questions pertain to (a) the types of assumptions needed to impose to identify these parameters, (b) how to estimate and make inferences about these parameters of interest once identification is established, (c) how to summarize treatment effect heterogeneity across doses to generate interpretable aggregated causal parameters, and whether traditional regression specifications based on TWFE recover a sensible and easy-to-understand causal parameter of interest. These questions are addressed in detail by Callaway2021b and deChaisemartin2022-continuous-treatment.
Callaway2021b highlights how, when no units are treated in period one, identification and estimation of $ATT(d|d)$’s (or their functionals) follows the binary case. They propose flexible nonparametric estimators for the $ATT(d|d)$ curve—the relationship between outcome changes (minus the average change for untreated units) and the dose $d$, making it possible to visualize and make inference about treatment effect heterogeneity across dosages. They also propose estimators that aggregate across dosage values and can be more precisely estimated. The identification of causal response parameters or ATE-type parameters, however, requires a stronger version of parallel trends that holds for potential outcomes at non-zero treatment doses. Under these strong parallel trends (and no-anticipation), they discuss estimation and inference procedures for the ACR curves and their summary measures. [^36] Interestingly, they also show that commonly used TWFE regression specifications are too rigid to lead to easy-to-interpret causal parameters of interest. In fact, they show that one can provide several different decompositions of the TWFE treatment coefficient depending on the specific causal parameter being used as a building block for the analysis, though every decomposition considered by them has some issues related to negative-weighting, additional “bias” terms, or non-interpretable weights that can distort inference. They emphasize that all this can be easily resolved by adopting the forward-engineering approach.
@@ A.3 Triple Differences
The causal interpretation of DiD estimates depends on the plausibility of their identification assumptions, which involve a no-anticipation and a parallel trends condition. In some applications, however, these assumptions may not hold, e.g., when the trends of average untreated outcomes among men and women vary across treatment groups. In these cases, a common empirical practice is to attempt to model these violations of parallel trends directly or to conduct sensitivity analysis (Freyaldenhoven_etal_2024; Rambachan2023). In some specific treatment designs in which treatment is rolled out to different units or groups (e.g., states), but is targeted to a specific subset (partition) of the population (e.g., women), it is possible to relax DiD-type parallel trends such that partition-specific and group-specific violations of parallel trends are allowed. Such setups are often referred to as Triple Differences (DDD). Since its introduction by Gruber1994, DDD has become very popular among empirical researchers—see Olden2022 for documentation. In this section, we provide a brief overview of the target parameters and identifying assumptions in DDD. We also highlight that, contrary to conventional wisdom, DDD procedures cannot generally be expressed as the difference between two DiD, especially when parallel trends assumptions are only plausible after conditioning on covariates or when treatment adoption is staggered. This discussion borrows heavily from Ortiz-Villavicencio_SantAnna_2025_DDD.
@@
When there are only two periods, $t=1,2$, two groups, $S\in\{2,\infty\}$, and covariates play no role in terms of identification, i.e., Assumption DDD-PT-GT-all holds without $X$ (or, equivalently, with $X=1$ for all units), Olden2022 show that one can identify $ATT(2,2)$ as the difference of two DiD estimands, i.e.,
@@
Ortiz-Villavicencio_SantAnna_2025_DDD show that DDD estimands cannot be written as the difference of two DiD estimands when covariates are important for identification (i.e., when parallel trends is only plausible after accounting for covariate-specific trends), or when treatment adoption is staggered over time and one wants to use not-yet-treated units as a comparison group (as is commonly done in DiD setups). They show how ignoring these considerations, and proceeding as if DDD was indeed just a difference of two DiDs, can lead to severely biased estimates for the $ATT(g,t)$’s. Ortiz-Villavicencio_SantAnna_2025_DDD also shows how one can avoid these issues by adopting a forward-engineering approach to the DDD problem. They propose regression-adjusted, inverse probability weighting, and doubly robust estimators for DDD setups that can reliably recover $ATT(g,t)$ and their associated summary parameters under mild assumptions. The paper discusses using multiple comparison groups to generate more precise estimates than simply using a single comparison group. Related, Strezhnev_DDD_2023 discusses several limitations of common two-way fixed effects regression specifications commonly used for DDD analysis.
Sometimes researchers use the term triple differences to mean different things, and often use different identification assumptions to estimate these different quantities. See, e.g., Caron2025 for a discussion on using a triple difference strategy to estimate treatment effect heterogeneity. We recommend that practitioners be transparent about target parameters, research designs, and identification assumptions to allow the research community to understand the goals and the differences between DDD procedures.
@@ A.4 Distributional DiD procedures
Our paper focuses on learning about average treatment effects in various DiD setups. However, approaches that embrace heterogeneity can also target quantities that describe heterogeneity other than average treatment effect parameters. In some settings, researchers may want more information about the distributional impacts of treatment participation. For instance, if a policymaker faces two different labor market programs with very similar average effects on earnings, they may prefer the one that potentially has a higher impact on the lower tail of the income distribution. Difference-in-Differences-type strategies can also be used to identify, estimate, and make inferences about various distributional features of the outcome of interest. This area has received a substantial amount of methodological consideration by econometricians in recent years; see, e.g., Athey2006, Bonhomme2011, Callaway2018, Callaway2019, Roth2023b, Ghanem2023_CiC, FernandezVal_2024_Dist_DiD, and references therein. For some empirical literature using distributional DiD procedures, see, e.g., Meyer_Viscusi_Durbin_1995, Finkelstein2008, and Cengiz2019, among many others.
An analysis of distributional quantities does not require different potential outcomes notation relative to Section 5.2; it just targets functionals of the potential outcome distributions other than their means. The first thing to notice is that there are several types of distributional causal parameters in the treated group that one may care about. The unique feature of them is that they are all functionals of $F_{Y_{t}(g)|G=g}(y)=\mathbb{P}(Y_{t}(g)\leq y|G=g)$ and $F_{Y_{t}(\infty)|G=g}(y)\equiv\mathbb{P}(Y_{t}(\infty)\leq y|G=g)$. Examples of such functionals include distributional treatment effects in time period $t$ among units first treated in period $g$ (denominated in probability units),
@@
where $F_{Y_{t}(g)|G=g}^{-1}(\tau)=inf\{y:F_{Y_{t}(g)|G=g}(y)\geq\tau\}$ denotes the $\tau$-quantile of $Y_{t}(g)$ among units in group $G=g$, and $F_{Y_{t}(\infty)|G=g}^{-1}(\tau)$ is defined analogously. Other functionals related to inequality measures can also be obtained; see, e.g., Firpo2016 for a discussion on this topic.
@@
We close this section by noting that there exist other types of distributional parameters of interest related to the distribution of the treatment effects in period $t$ among the units in group $g$, $\mathbb{P}(Y_{t}(g)-Y_{t}(\infty)\leq y|G=g)$. In general, such causal quantities cannot be point-identified, as discussed in Heckman1997b, FanYu2012 and Callaway2021c. However, often one can still partially identify such policy-relevant parameters under different restrictions. We refer the reader to Callaway2021c for a more detailed discussion of this topic.
@@ A.5 Repeated cross-sections and unbalanced panel data
An appealing feature of DiD procedures is that, although helpful, a balanced panel is not a requirement for DiD analyses, which can also be deployed with repeated cross-sectional data or unbalanced panels. Indeed, as discussed in Section 3.3 and made explicit in equation 3.6, the $2\times 2$ building block in unconditional DiD analyses only involves averages that are group and time specific, and does not require the same unit to be observed in all periods. As discussed in Callaway2021, the same applies to unconditional staggered adoption setups, and one need not enforce a balanced panel even within each subset of the data used to estimate the $ATT(g,t)$ building blocks. One caveat is that the interpretation of the parameter of interest may change, which we discuss more below.
@@
where $T_{i,t}$ is an indicator if unit $i$ is observed in period $t$, $Y_{i}(d)=T_{i,t=2}\leavevmode\nobreak\ Y_{i,t=2}(d)+T_{i,t=1}\leavevmode\nobreak\ Y_{i,t=1}(d)$ is the potential outcome for unit $i$, $D_{i}=1\{G_{i}=2\}$ is a treatment group dummy that equals one if a unit is first treated in period two and zero if it is untreated in both periods. We also set $X_{i}$ to be a vector of (pre-treatment) covariates. Note that even here we already use the stationarity condition that the joint distribution of $(D_{i},X_{i})$ is invariant to $T_{i,t=2}$ to move from the first to the second line and establish (A.1).
To identify $ATT(2)$ it is often constructive to first establish the identification of its conditional-on-covariates analog, i.e., the conditional ATT in period two among units with covariates $X_{i}$, $ATT_{X_{i}}(2)$. This is exactly how we proceeded in Section 4.2. Under the stationarity condition, and similarly to (A.1), we can express this quantity as [^38] To guarantee that all the conditional expectations in (A.2) are well-defined, we need an overlap condition that guarantees that $P(T_{i,t=2}=1,D_{i}=1|X_{i})>0$ We discuss this below.
@@
Next, we have to establish the identification of this quantity. As expected, we will again use conditional parallel trends, no-anticipation, and overlap assumptions. The no-anticipation condition used here is the same as in the main text, i.e., we assume Assumption NA. The conditional parallel trends and overlap assumptions need to be modified, as we now work with multiple partitions of the data depending on treatment status and the period a unit is observed. In this sense, we modify Assumptions CPT and SO to the following related, but different, assumptions. These modifications are warranted regardless of whether compositional changes are present; this step is instead tied to data structure. [^39] In setups where we rule out compositional changes and impose the stationarity condition that the joint distribution of $(D_{i},X_{i})$ is invariant to $T_{i,t=2}$, we may not need to modify Assumption CPT. We do it here for transparency purposes.
@@
With this modification, we can now show that when Assumptions NA, CPT-RCS, and SO hold, the conditional ATT parameter $ATT_{X_{i}}(2)$ is identified by [^40] We also require the assumption that the pooled repeated cross-section data $\left\{Y_{i},D_{i},X_{i},T_{i,t=2},T_{i,t=1}\right\}_{i=1}^{n}$ is $iid$, though this is fairly standard and uncontroversial; see, e.g., Abadie2005 and SantAnna2020. We maintain this condition as an assumption throughout this section.
@@
This step provides a methodological justification to estimate the $ATT_{X_{i}}(2)$’s using four conditional expectations that only use the available data. In addition, it highlights that, under the stationarity assumption, once we learn the $ATT_{X_{i}}(2)$s, we can aggregate them using the covariate distribution of treated units available from both time periods to get the $ATT(2)$. More formally, $ATT(2)$ is identified by
@@
where $ATT_{X_{i}}(2)$ is given by (A.4). This is the second place where ruling out compositional changes and imposing the stationarity assumption helps: it allows you to use covariates from treated units in the entire dataset to recover $ATT(2)$. The fact that you can pool data across all periods to learn about $ATT(2)$ translates to gains in power, as formally discussed by SantAnna2020 and Santanna_Qi_DiD_compositional_2023. The third place where the stationarity condition affects the analysis is in the characterization of how “the most precise” (regular and asymptotically linear) estimator for the $ATT(2)$ should look. This point relates to the semi-parametric efficiency bound and the construction of efficient (and doubly robust) estimators. As these points are slightly more technical, we refer the readers to SantAnna2020 and Santanna_Qi_DiD_compositional_2023 for more details about this.
Overall, ruling out compositional changes over time allows for the pooling of information across the entire dataset, which has an impact on the definition of target parameters and leads to more precise inference procedures. But what happens when this condition is not plausible? How does this affect the analysis?
First, this matters for the definition of the treatment effect of interest. In setups where the sampling varies across periods, we do not have a single notion of $ATT(2)$. Instead, we need to accommodate the fact that the $ATT(2)$ may vary across units sampled from different periods. As such, when we do not rule out compositional changes, we must be explicit about the treated subpopulation that we are interested in. It is common to focus on the average treatment effect in period two among treated units that are also sampled in period two, that is, [^41] One may also be interested in the ATT in period two among treated units that are sampled in period one. The arguments required to establish (point) identification of this parameter differ from those we use here. A main challenge is that we do not observe $\mathop{}\!\mathbb{E}[Y_{i,t=2}(1)|D_{i}=1,T_{i,t=1}=1]$, and the parallel trends assumption we leverage does not involve treated potential outcomes.
@@
Although (A.5) has the same statistical estimand as (A.1)—i.e., the formulas on the right-hand side of the equation coincide—it has a very different interpretation. It is the $ATT(2)$ among units sampled in period $2$, and is not an “overall” $ATT(2)$. One may think this difference is merely cosmetic, but as discussed below, this has implications for constructing estimands when covariates are important for identification. Where covariates do not play an important role, it is simply a matter of changing the interpretation of your reported estimates (which also applies to unconditional staggered setups, to be clear).
When covariates do play an important role, however, and when Assumptions CPT-RCS, NA and SO-RCS hold—the conditional ATT parameter among units sampled in period two, $ATT_{X_{i}}(2|T_{t=2})$ is identified by
@@
We close this section by highlighting that, in practice, it is possible to test for compositional changes by comparing the estimates from estimators that impose it and those that do not. Santanna_Qi_DiD_compositional_2023 discuss Hausman-type tests in the two-period setting, though one can extend those to more general setups. We also highlight that when it comes to DiD setups with staggered adoption, some equivalence results discussed in Section 5.2 no longer hold with repeated cross-sections or unbalanced panel data. For instance, the Sun2021 regression-based strategy to estimate $ATT(g,t)$’s using (5.11) no longer coincide with Callaway2021 estimators using the analogous of (5.10), i.e.,
@@
with $\overline{Y}_{G=a,t=s}$ being the sample mean of $Y$ among units that belong in group $G=a$ and are observed in period $t=s$. In fact, it is unclear exactly what estimand is being recovered when one uses (5.11) with an unbalanced panel. If one replaces unit fixed effects with treatment group dummies in (5.11), such equivalence is restored, though we suspect that many practitioners do not use this alternative specification. In general, we caution against extrapolating from a well-motivated regression specification that was studied under one specific setup to another related but inherited different framework. This practice has led to many issues in DiD, which can be fully avoided by adopting the forward-engineering approach discussed in this paper.

### v2

@@ A.2 DiD setups with continuous or multi-valued treatments
Our paper focuses on binary treatments, but many treatments take multiple values or are even continuous. A number of recent papers have studied this particular type of treatment design. These include Callaway2021b; CallawayGoodmanBaconSantAnna2024; deChaisemartin2022-continuous-treatment; and deChaisemartinDHaultfoeuilleVazquezBare2024. Here we focus on a two-period setting in which no unit is treated in period one and some units receive a treatment with varying intensities (or doses) in period two. Most of the key results that distinguish multi-valued from binary treatments are evident with two periods (CallawayGoodmanBaconSantAnna2024).
We now need to define potential outcomes that reflect varying treatment intensity. We denote $Y_{i,t}(0,d)$ as potential outcomes for unit $i$ in period $t$ if they are untreated in period one and receive treatment dosage $d$ in period two. As we focus on setups where all units are untreated in period one, we simplify notation and index potential outcomes by treatment intensity in period two; that is, $Y_{it}(d)=Y_{i,t}(0,d)$. An important feature is that $d$ is not restricted to $\{0,1\}$ and can take on richer treatment intensities instead. We denote the treatment dosage for unit $i$ as $D_{i}$ in period two and stress that in this context, our notion of the treatment group is tied to units’ treatment dosage: groups are defined by their treatment dosage in period 2.
@@
reflect the average effect of dose $d$ relative to no treatment. Here $ATT(d|d^{\prime})$ is the average treatment effect for units that experienced dose $d^{\prime}$; when $d^{\prime}=d$, it is the $ATT$ among units that received dose $d$. On the right side, $ATE(d)$ is defined analogously, except that it is the effect on the overall population. Of course, one can also aggregate these dose-specific parameters to form more precisely estimable summary quantities; see, e.g., Callaway2021b.
The two treatment effect parameters above provide average treatment effects in levels, and so one reason why they vary could be because $d$ itself varies. To account for the differences in $d$, one may be interested in “per-dosage” effects:
@@
Here $ACRT(d|d)$ equals the derivative of the average potential outcome in period two for units that received dose $d$ evaluated at $d$—this is equivalent to the derivative of $ATT(l|d)$ with respect to $l$, evaluated at $l=d$. We can interpret $ACR(d)$ analogously. [^34] For discrete treatments, ACR’s are defined in a similar way but with a slightly different notation to accommodate the discreteness of $d$: $ACRT(d_{j}|d_{k})=\mathop{}\!\mathbb{E}[Y_{t=2}(d_{j})-Y_{t=2}(d_{j-1})|D=d_{k}]$, and $ACR(d_{j})=\mathop{}\!\mathbb{E}[Y_{t=2}(d_{j})-Y_{t=2}(d_{j-1})]$.
The relevant questions pertain to (a) what assumptions are needed to impose to identify these parameters, (b) how to estimate and make inferences about these parameters of interest once identification is established, (c) how to summarize treatment effect heterogeneity across doses to generate interpretable aggregated causal parameters, and (d) whether traditional regression specifications based on TWFE recover a sensible and easy-to-understand causal parameter of interest. These questions are addressed in detail by Callaway2021b and deChaisemartin2022-continuous-treatment.
Callaway2021b highlight how, when no units are treated in period one, identification and estimation of $ATT(d|d)$’s (or their functionals) follows the binary case. They propose flexible nonparametric estimators for the $ATT(d|d)$ curve—the relationship between outcome changes (minus the average change for untreated units) and the dose $d$, making it possible to visualize and make inference about treatment effect heterogeneity across dosages. They also propose estimators that aggregate across dosage values and can be more precisely estimated. The identification of causal response parameters or ATE-type parameters, however, requires a stronger version of parallel trends that holds for potential outcomes at non-zero treatment doses. Under these strong parallel trends and no anticipation assumptions, they discuss estimation and inference procedures for the ACR curves and their summary measures. [^35] Interestingly, they also show that commonly used TWFE regression specifications are too rigid to lead to easy-to-interpret causal parameters of interest. In fact, they show that one can provide several different decompositions of the TWFE treatment coefficient depending on the specific causal parameter being used as a building block for the analysis, though every decomposition considered by them has some issues related to negative-weighting, additional “bias” terms, or non-interpretable weights that can distort inference. They emphasize that all this can be easily resolved by adopting the forward-engineering approach.
@@ A.3 Triple differences
The causal interpretation of DiD estimates depends on the plausibility of their identification assumptions, which involve a no-anticipation and a parallel trends condition. In some applications, however, these assumptions may not hold—for example, when the trends of average untreated outcomes among men and women vary across treatment groups. In these cases, a common empirical practice is to attempt to model these violations of parallel trends directly or to conduct sensitivity analysis (Freyaldenhoven_etal_2024; Rambachan2023). In some specific treatment designs in which treatment is rolled out to different units or groups (e.g., states), but is targeted to a specific subset (partition) of the population (e.g., women), it is possible to relax DiD-type parallel trends so that partition-specific and group-specific violations of parallel trends are allowed. Such setups are often referred to as “triple differences” (DDD). Since its introduction by Gruber1994, DDD has become very popular among empirical researchers—see Olden2022 for documentation. In this section, we provide a brief overview of the target parameters and identifying assumptions in DDD. We also highlight that, contrary to conventional wisdom, DDD procedures cannot generally be expressed as the difference between two DiD, especially when parallel trends assumptions only hold after conditioning on covariates or when treatment adoption is staggered. This discussion borrows heavily from Ortiz-Villavicencio_SantAnna_2025_DDD.
@@
When there are only two periods, $t=1,2$, and two groups, $S\in\{2,\infty\}$, and covariates play no role in terms of identification—that is, Assumption DDD-PT-GT-all holds without $X$ (or, equivalently, with $X=1$ for all units)—Olden2022 show that one can identify $ATT(2,2)$ as the difference of two DiD estimands:
@@
Ortiz-Villavicencio_SantAnna_2025_DDD show that DDD estimands cannot be written as the difference of two DiD estimands when covariates are important for identification, or when treatment adoption is staggered over time and one wants to use not-yet-treated units as a comparison group (as is commonly done in DiD setups). They show how ignoring these considerations and proceeding as if DDD were indeed just a difference of two DiDs can lead to severely biased estimates for the $ATT(g,t)$’s. Ortiz-Villavicencio_SantAnna_2025_DDD also show how one can avoid these issues by adopting a forward-engineering approach to the DDD problem. They propose regression-adjusted, inverse probability weighting, and doubly robust estimators for DDD setups that can reliably recover $ATT(g,t)$ and their associated summary parameters under mild assumptions. The paper discusses using multiple comparison groups to generate more precise estimates than simply using a single comparison group. Relatedly, Strezhnev_DDD_2023 discusses several limitations of common two-way fixed effects regression specifications commonly used for DDD analysis.
Sometimes researchers use the term “triple differences” to mean different things and often use different identification assumptions to estimate these different quantities. Caron2025 discusses using a triple difference strategy to estimate treatment effect heterogeneity. We recommend that practitioners be transparent about target parameters, research designs, and identification assumptions to allow the research community to understand the goals and the differences between DDD procedures.
@@ A.4 Distributional DiD procedures
Our paper focuses on learning about average treatment effects in various DiD setups. However, approaches that embrace heterogeneity can also target quantities that describe heterogeneity other than average treatment effect parameters. In some settings, researchers may want more information about the distributional impacts of treatment participation. For instance, if a policymaker faces two different labor market programs with very similar average effects on earnings, they may prefer the one that potentially has a higher impact on the lower tail of the income distribution. Difference-in-Differences-type strategies can also be used to identify, estimate, and make inferences about various distributional features of the outcome of interest. This area has received a substantial amount of methodological consideration by econometricians in recent years; see Athey2006, Bonhomme2011, Callaway2018, Callaway2019, Roth2023b, Ghanem2023_CiC, FernandezVal_2024_Dist_DiD, and references therein. For some empirical literature using distributional DiD procedures, see Meyer_Viscusi_Durbin_1995, Finkelstein2008, and Cengiz2019, among many others.
An analysis of distributional quantities does not require different potential outcomes notation from Section 5.2; it just targets functionals of the potential outcome distributions other than their means. The first thing to notice is that there are several types of distributional causal parameters in the treated group that one may care about. The unique feature of them is that they are all functionals of $F_{Y_{t}(g)|G=g}(y)=\mathbb{P}(Y_{t}(g)\leq y|G=g)$ and $F_{Y_{t}(\infty)|G=g}(y)\equiv\mathbb{P}(Y_{t}(\infty)\leq y|G=g)$. Examples of such functionals include distributional treatment effects in time period $t$ among units first treated in period $g$ (denominated in probability units),
@@
where $F_{Y_{t}(g)|G=g}^{-1}(\tau)=inf\{y:F_{Y_{t}(g)|G=g}(y)\geq\tau\}$ denotes the $\tau$-quantile of $Y_{t}(g)$ among units in group $G=g$, and $F_{Y_{t}(\infty)|G=g}^{-1}(\tau)$ is defined analogously. Other functionals related to inequality measures can also be obtained; see Firpo2016 for a discussion on this topic.
@@
We close this section by noting that there exist other types of distributional parameters of interest related to the distribution of the treatment effects in period $t$ among the units in group $g$, $\mathbb{P}(Y_{t}(g)-Y_{t}(\infty)\leq y|G=g)$. In general, such causal quantities cannot be point identified, as discussed in Heckman1997b, FanYu2012 and Callaway2021c. However, often one can still partially identify such policy-relevant parameters under different restrictions. We refer the reader to Callaway2021c for a more detailed discussion of this topic.
@@ A.5 Repeated cross-sections and unbalanced panel data
An appealing feature of DiD procedures is that, although helpful, a balanced panel is not a requirement for DiD analyses, which can also be deployed with repeated cross-sectional data or unbalanced panels. Indeed, as discussed in Section 3.3 and made explicit in equation 3.6, the $2\times 2$ building block in unconditional DiD analyses involves only averages that are group and time specific, and does not require the same unit to be observed in all periods. As discussed in Callaway2021, the same applies to unconditional staggered adoption setups, and one need not enforce a balanced panel even within each subset of the data used to estimate the $ATT(g,t)$ building blocks. One caveat is that the interpretation of the parameter of interest may change, which we discuss more below.
@@
where $T_{i,t}$ is an indicator if unit $i$ is observed in period $t$, $Y_{i}(d)=T_{i,t=2}\leavevmode\nobreak\ Y_{i,t=2}(d)+T_{i,t=1}\leavevmode\nobreak\ Y_{i,t=1}(d)$ is the potential outcome for unit $i$, $D_{i}=1\{G_{i}=2\}$ is a treatment group dummy that equals one if a unit is first treated in period two and zero if it is untreated in both periods. We also set $X_{i}$ to be a vector of (pre-treatment) covariates. Note that even here, we already use the stationarity condition that the joint distribution of $(D_{i},X_{i})$ is invariant to $T_{i,t=2}$ to move from the first to the second line and establish (A.1).
To identify $ATT(2)$ it is often constructive to first establish the identification of its conditional-on-covariates analog; that is, the conditional ATT in period two among units with covariates $X_{i}$, $ATT_{X_{i}}(2)$. This is exactly how we proceeded in Section 4.2. Under the stationarity condition, and similarly to (A.1), we can express this quantity as [^37] To guarantee that all the conditional expectations in (A.2) are well-defined, we need an overlap condition that guarantees that $P(T_{i,t=2}=1,D_{i}=1|X_{i})>0$. We discuss this below.
@@
Next, we have to establish the identification of this quantity. As expected, we will again use conditional parallel trends, no-anticipation, and overlap assumptions. The no-anticipation condition used here is the same as the one in the main text. The conditional parallel trends and overlap assumptions need to be modified, as we now work with multiple partitions of the data depending on treatment status and the period a unit is observed. In this sense, we modify Assumptions CPT and SO to the following related, but different, assumptions. These modifications are warranted regardless of whether compositional changes are present; this step is instead tied to data structure. [^38] In setups where we rule out compositional changes and impose the stationarity condition that the joint distribution of $(D_{i},X_{i})$ is invariant to $T_{i,t=2}$, we may not need to modify Assumption CPT. We do it here for transparency purposes.
@@
With this modification, we can now show that when Assumptions NA, CPT-RCS, and SO hold, the conditional ATT parameter $ATT_{X_{i}}(2)$ is identified by [^39] We also require the assumption that the pooled repeated cross-section data $\left\{Y_{i},D_{i},X_{i},T_{i,t=2},T_{i,t=1}\right\}_{i=1}^{n}$ is $iid$, though this is fairly standard and uncontroversial; see, for instance, Abadie2005 and SantAnna2020. We maintain this condition as an assumption throughout this section.
@@
This step provides a methodological justification to estimate the $ATT_{X_{i}}(2)$’s using four conditional expectations that use only the available data. In addition, it highlights that, under the stationarity assumption, once we learn the $ATT_{X_{i}}(2)$’s, we can aggregate them using the covariate distribution of treated units available from both time periods to get the $ATT(2)$. More formally, $ATT(2)$ is identified by
@@
where $ATT_{X_{i}}(2)$ is given by (A.4). This is the second place where ruling out compositional changes and imposing the stationarity assumption helps: it allows you to use covariates from treated units in the entire dataset to recover $ATT(2)$. The fact that you can pool data across all periods to learn about $ATT(2)$ translates to gains in power, as formally discussed by SantAnna2020 and Santanna_Qi_DiD_compositional_2023. The third place where the stationarity condition affects the analysis is in the characterization of how “the most precise” (regular and asymptotically linear) estimator for the $ATT(2)$ should look. This point relates to the semi-parametric efficiency bound and the construction of efficient (and doubly robust) estimators. As these points are slightly more technical, we refer the readers to SantAnna2020 and Santanna_Qi_DiD_compositional_2023 for more details.
Overall, ruling out compositional changes over time allows for the pooling of information across the entire dataset, which has an impact on the definition of target parameters and leads to more precise inference procedures. But what happens when this condition fails? How does this affect the analysis?
First, this matters for the definition of the treatment effect of interest. In setups where the sampling varies across periods, we do not have a single notion of $ATT(2)$. Instead, we need to accommodate the fact that the $ATT(2)$ may vary across units sampled from different periods. Thus, when we do not rule out compositional changes, we must be explicit about the treated subpopulation that we are interested in. It is common to focus on the average treatment effect in period two among treated units that are also sampled in period two, that is, [^40] One may also be interested in the ATT in period two among treated units that are sampled in period one. The arguments required to establish (point) identification of this parameter differ from those we use here. A main challenge is that we do not observe $\mathop{}\!\mathbb{E}[Y_{i,t=2}(1)|D_{i}=1,T_{i,t=1}=1]$, and the parallel trends assumption we leverage does not involve treated potential outcomes.
@@
Although (A.5) has the same statistical estimand as (A.1)—that is, the formulas on the right-hand side of the equation coincide—it has a very different interpretation. It is the $ATT(2)$ among units sampled in period $2$, and is not an “overall” $ATT(2)$. One may think this difference is merely cosmetic, but as discussed below, this has implications for constructing estimands when covariates are important for identification. Where covariates do not play an important role, it is simply a matter of changing the interpretation of your reported estimates (which also applies to unconditional staggered setups, to be clear).
When covariates do play an important role, however, and when Assumptions CPT-RCS, NA and SO-RCS hold, the conditional ATT parameter among units sampled in period two, $ATT_{X_{i}}(2|T_{t=2})$ is identified by
@@
We close this section by highlighting that, in practice, it is possible to test for compositional changes by comparing the estimates from estimators that impose it and those that do not. Santanna_Qi_DiD_compositional_2023 discuss Hausman-type tests in the two-period setting, though one can extend those to more general setups. We also highlight that when it comes to DiD setups with staggered adoption, some equivalence results discussed in Section 5.2 no longer hold with repeated cross-sections or unbalanced panel data. For instance, the Sun2021 regression-based strategy to estimate $ATT(g,t)$’s using (5.11) no longer coincides with Callaway2021’s (Callaway2021) estimators using the analog of (5.10):
@@
where $\overline{Y}_{G=a,t=s}$ is the sample mean of $Y$ among units that belong in group $G=a$ and are observed in period $t=s$. In fact, it is unclear exactly what estimand is being recovered when one uses (5.11) with an unbalanced panel. If one replaces unit fixed effects with treatment group dummies in (5.11), such equivalence is restored, though we suspect that many practitioners do not use this alternative specification. In general, we caution against extrapolating from a well-motivated regression specification that was studied under one specific setup to another related but inherited different framework. This practice has led to many issues in DiD, which can be fully avoided by adopting the forward-engineering approach discussed in this paper.